OpenStudy (anonymous):

The average speed of a Greyhound bus from Lansing to Detroit is 105.5 km/h. On the return trip from Detroit to Lansing the average speed is 55.2 km/h on the same road due to heavy traffic. What is the average speed of the bus for the round trip?

jimthompson5910 (jim_thompson5910):

We're going to use the equation d = rt d = distance r = rate (ie speed) t = time

jimthompson5910 (jim_thompson5910):

When it says "The average speed of a Greyhound bus from Lansing to Detroit is 105.5 km/h" this means d = rt d = 105.5t now solve for t to get t = d/105.5

jimthompson5910 (jim_thompson5910):

Then it says "On the return trip from Detroit to Lansing the average speed is 55.2 km/h on the same road due to heavy traffic." so d is still the same (same distance), but the time is now different. So let's say the new time is x (instead of t d = r*t d = 55.2*x solve for this new time x to get x = d/55.2

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

yes! sorry I just saw this!

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

The time it takes to go from Lansing to Detroit is d/105.5 hours The time it takes to come back is d/55.2 hours

jimthompson5910 (jim_thompson5910):

So the total time is just the sum of the two times, ie add them up d/105.5+d/55.2 (d/105.5)*(55.2/55.2)+d/55.2 (55.2d)/5823.6+d/55.2 (55.2d)/5823.6+(d/55.2)*(105.5/105.5) (55.2d)/5823.6+(105.5d)/5823.6 (55.2d+105.5d)/5823.6 160.7d/5823.6 The total time, in terms of the distance d, is 160.7d/5823.6 hours This is a lot to take in, so go over it a few times and let me know when you're ready to move on.

OpenStudy (anonymous):

Okay I got it

OpenStudy (anonymous):

But where does the 55.2/55.2 come from?

OpenStudy (anonymous):

Oh wait, okay I see

jimthompson5910 (jim_thompson5910):

in order to add the fractions, the denominators had to be the same so that's how you fix that

jimthompson5910 (jim_thompson5910):

once you get it to (55.2d)/5823.6+(105.5d)/5823.6, you can add the fractions

OpenStudy (anonymous):

so 160.7/5823.6?

jimthompson5910 (jim_thompson5910):

the distance d is in there too

jimthompson5910 (jim_thompson5910):

so the total time is 160.7d/5823.6

jimthompson5910 (jim_thompson5910):

if d wasn't in there, then yeah, it would be 160.7/5823.6

OpenStudy (anonymous):

Then do you solve for d?

jimthompson5910 (jim_thompson5910):

so we know the total distance is d+d = 2d (basically double the distance to cover going there and back) and the total time is 160.7d/5823.6

jimthompson5910 (jim_thompson5910):

average speed = (total distance)/(total time)

jimthompson5910 (jim_thompson5910):

if you divide the two, you'll find that the 'd' variables divide and cancel out you'll be left with a number without variables

OpenStudy (anonymous):

so I get .02759, but thats wrong when I put it in...did I do that wrong?

jimthompson5910 (jim_thompson5910):

you should have average speed = (2d)/(160.7d/5823.6) simplify that

jimthompson5910 (jim_thompson5910):

the answer you got is way too small

OpenStudy (anonymous):

72.48?

jimthompson5910 (jim_thompson5910):

I'm getting 72.47790916, but rounding that to 2 decimal places gives 72.48

OpenStudy (anonymous):

Okay, thank you so much!

jimthompson5910 (jim_thompson5910):

sure thing